Two trains going in opposite directions leave at the same time. Train B travels 5 mph faster than train A. In 6 hours the trains are 810 miles apart. Find the speed of each.
Answer by stanbon(75887)

Two trains going in opposite directions leave at the same time. Train B travels 5 mph faster than train A. In 6 hours the trains are 810 miles apart. Find the speed of each.
---
Let slower train speed be x.
The faster train speed is x+5.
---
Combined separation speed is 2x+5
---
Distance = Distance
6(2x+5) = 810
2x+5 = 135
2x = 130
x = 65 mph (slower train rate)
x+5 = 70 mph (faster train rate)
===================
Cheers,
Stan H.
=========